# Time Travel in LLMs: Tracing Data Contamination in Large Language Models

### v1

@@ 6 Conclusion
We proposed a methodology to detect data contamination in LLMs, assuming the lack of access to their training data. Our approach began by pinpointing data contamination at the instance level. This was achieved by prompting the LLM to produce the replica of the secondary segment of a dataset instance given its initial random segment, dataset name, and partition type, a process we called “guided instruction.” From here, we adopted a set of rules to generalize from instance-level to broader partition-level contamination. This involved leveraging statistically significant differences from BLEURT and/or ROUGE-L scores between guided and general instructions, as well as evaluations from GPT-4 using in-context learning prompting.
Our evaluation spanned 28 different settings, including seven datasets along with their respective train and test/validation partitions for both GPT-4 and GPT-3.5. Our findings indicated that while the replication technique via guided instruction is notably effective, the most accurate approach that was closely aligned with human judgments for detecting data contamination was the few-shot in-context learning prompt with GPT-4, integrated a few exemplary instances from human assessment in the input prompt. This method yielded a success rate in pinpointing data contamination across 14/14 scenarios for GPT-4 and 13/14 for GPT-3.5.
Appendix A A Complete List of All Guided and General Instructions
Table 3 presents a thorough collection of all the guided and general instructions employed throughout our study. These instructions utilize placeholders, each of which is replaced with pertinent information as follows: {dataset_name} is substituted with the name of the dataset; {split_name} is replaced with the title of the split; {input} stands in for the preliminary part of the dataset instance cut at the tail randomly, or in the case of NLI-based datasets, it encompasses the whole initial sentence; and {label} is exchanged with the label corresponding to the incomplete input instance.
Table 3: A comprehensive list of all guided and general instructions used in our experiments.
@@ Appendix A A Complete List of All Guided and General Instructions
Appendix B Statistical Analysis: Bootstrap Resampling
We examine the statistical significance of results stemming from guided versus general instructions. A bootstrap resampling technique, involving 10,000 samples in the resampling process, is employed for this investigation (Efron 1979; Efron and Tibshirani 1993; Efron 2003). We concentrate on the alternative hypothesis that posits guided instructions produce outcomes closer to reference instances than those generated from general instructions, as evaluated by fluency, quality, and similarity. The performance metrics utilized here are BLEURT and ROUGE-L scores. We regard the ROUGE-L and BLEURT scores as statistically significant if the $p$-values $\leq 0.05$. We highlight the statistically significant results by marking them with an asterisk in Table 2.
Appendix C Few-shot In-context Learning Prompt
Figure 2 showcases the few-shot ICL prompt employed to evaluate the model-generated candidate against the reference text using GPT-4. Within this prompt, we present GPT-4 with one precise match and four exemplary instances of near-exact matches, all pre-labeled by human evaluation. These examples guide GPT-4 in discerning the difference between near-exact and inexact matches, in line with human assessment.
@@ Appendix C Few-shot In-context Learning Prompt
Figure 2: A display of the few-shot ICL prompt utilized for instance-level data contamination detection using GPT-4. In this illustration, examples 1 through 4 remain constant within the prompt, while example 5 is updated with a new input reference and candidate for evaluation, depending on whether there is an exact, nearly exact, or inexact match.
Appendix D Illustrations of Exact, Near-exact, and Inexact Matches
@@
Displayed in Table 4 are examples of exact, near-exact, and inexact replicas of the reference instance when guided instruction and GPT-4 are used. This table also includes computed metrics such as ROUGE-L, BLEURT, and results from human and GPT-4 evaluations. In addition, Table 5 showcases comparative outcomes for the same examples using general instruction.
@@
Table 4: Examples of exact, near-exact, and inexact matches along with their respective BLEURT and ROUGE-L scores, and judgments from GPT-4 ICL and human evaluations. These examples are generated by GPT-4, as the underlying language model.
Match Types and Metrics Reference Instance and Its Replica by Guided Instruction
@@
Reference Instance:Bromwell High is a cartoon comedy. It ran at the same time as some other programs about school life,such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’ssatire is much closer to reality than is ”Teachers”. The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ……… at ………. High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!Label: 1 (positive)
@@
BLEURT Score: 0.99 ROUGE-L Score: 1.00 GPT-4 Eval.: Exact Match Human Eval.: Exact Match Instruction: You are provided with the first piece of an instance from the train split of the IMDB dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: 1 (positive) First Piece: Bromwell High is a cartoon comedy. It ran at the same time as some other programs about school life, such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’s satire is much closer to reality than is ”Teachers”. Second Piece: The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ......... at .......... High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!
@@
BLEURT Score: 0.87 ROUGE-L Score: 0.82 GPT-4 Eval.: Near-exact match Human Eval.: Near-exact match Instruction: You are provided with Sentence 1 from the train split of the RTE dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: Although they were born on different planets, Oscar-winning actor Nicolas Cage’s new son and Superman have something in common - both were named Kal-el. Label: 1 (entailment) Sentence 2: Nicolas Cage’s new son is named Kal-el.
Reference Instance:Kim is about to tell mom that Harry bought a new sofa, and he needs grey pillows.
BLEURT Score: 0.48 ROUGE-L Score: 0.12 GPT-4 Eval.: Inexact match Human Eval.: Inexact match Instruction: You are provided with the first piece of a summary from the test split of the SAMSum dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: Kim is about to tell mom that Harry bought Second Piece: a new car but is worried mom will be upset. Kim is advised to tell mom in a positive way, focusing on Harry’s happiness.
@@
Metrics Reference Instance and Its Replica by General Instruction
Reference Instance:Bromwell High is a cartoon comedy. It ran at the same time as some other programs about school life,such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’ssatire is much closer to reality than is ”Teachers”. The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ……… at ………. High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!Label: 1 (positive)
@@
Reference Instance:Kim is about to tell mom that Harry bought a new sofa, and he needs grey pillows.
@@
Appendix E Detailed Description of Datasets
@@
Recognizing Textual Entailment (RTE) Dataset. The Recognizing Textual Entailment (RTE) dataset originates from a succession of annual textual entailment challenges. These datasets were combined by the authors of the benchmark using data from four different editions: RTE1 (Dagan, Glickman, and Magnini 2005), RTE2 (Haim et al. 2006), RTE3 (Giampiccolo et al. 2007), and RTE5 (Bentivogli et al. 2009). The examples within these datasets were primarily formulated using text from news and Wikipedia sources. To maintain consistency, all these datasets were adapted into a two-class split. For those datasets that initially consisted of three classes, the categories of ”neutral” and ”contradiction” were combined to form a single class termed ”not entailment”. The RTE dataset combined has 2,490 examples for training, 277 examples for validation, and 3,000 examples for testing.
Winograd Natural Language Inference (WNLI) Dataset. The WNLI (Winograd Natural Language Inference) dataset is a benchmark for natural language understanding tasks, particularly for evaluating coreference resolution and pronoun disambiguation in context. The dataset is derived from the original Winograd Schema Challenge (Levesque, Davis, and Morgenstern 2012) and contains sentence pairs where a pronoun needs to be resolved by determining whether it refers to the same entity as the previous sentence. While the dataset has a balanced training set between two classes, the test set is imbalanced, with 635 training examples, 146 testing examples, and 71 validation examples.

### v2

@@ 6 Conclusion
We proposed a novel method to detect data contamination in LLMs, assuming no access to their pre-training data. Our approach begins by pinpointing data contamination at the instance level. This was achieved by prompting the LLM to produce the replica of the secondary segment of a dataset instance given its random-length initial segment, dataset name, and partition type, a process we called “guided instruction.” From here, we adopted a set of rules to generalize from instance-level to broader partition-level contamination. This involved leveraging statistically significant differences from BLEURT and ROUGE-L scores between generated completions by guided and general instructions, as well as evaluations from GPT-4 with few-shot in-context learning prompting.
Our evaluation spanned 28 different settings, including seven datasets along with their respective train and test/validation partitions and two LLMs: GPT-4 and GPT-3.5. Our findings indicated that while the replication technique via guided instruction is notably effective, the most accurate evaluation approach that was closely aligned with human judgments for detecting data contamination was the few-shot in-context learning prompt with GPT-4, which integrates a few example instances from human assessments in the input prompt. This method yielded a success rate in pinpointing data contamination across 14/14 scenarios for GPT-4 and 13/14 for GPT-3.5.
Appendices
Appendix A List of All Guided and General Instructions
Table 5 presents a thorough collection of all the guided and general instructions employed throughout our study.
Table 5: A comprehensive list of all guided and general instructions used in our experiments. Placeholders include: {split_name} for the partition (or split) name; {dataset_name} for the dataset name; {input} for the first part of the dataset instance cut at the tail randomly or the whole first sentence in NLI-based datasets; and {label} for the corresponding label of the incomplete input instance.
@@ Appendix A List of All Guided and General Instructions
Appendix B Few-shot In-context Learning Prompt
Figure 3 showcases the few-shot ICL prompt employed to evaluate the model-generated candidate against the reference text using GPT-4. Within this prompt, we present GPT-4 with one exact match and three exemplary instances of near-exact matches, all pre-labeled by human evaluation. These examples guide GPT-4 in discerning the difference between near-exact and inexact matches, in line with human assessment.
@@ Appendix B Few-shot In-context Learning Prompt
Figure 3: A display of the few-shot ICL prompt utilized for instance-level data contamination detection using GPT-4. In this illustration, examples 1 through 4 are part of the prompt, while example 5 is updated with a new input reference and candidate for evaluation, depending on whether there is an exact, near-exact, or inexact match. While Example 1 represents an exact match, the other examples display variations indicating near-exact matches: Example 2 reveals a scenario where the candidate text has substantial overlap with the reference but includes added details; Examples 3 and 4 highlight situations where the candidate text possesses both semantic and structural similarity to the reference text.
Appendix C Illustrations of Exact, Near-exact, and Inexact Matches
@@
Displayed in Table 6 are examples of exact, near-exact, and inexact replicas of the reference instance when guided instruction and GPT-4 are used. This table also includes computed metrics such as ROUGE-L, BLEURT, and results from human and GPT-4 few-shot ICL evaluations. In addition, Table 7 showcases comparative outcomes for the same examples using general instruction.
@@
Table 6: Examples of exact, near-exact, and inexact matches along with their respective BLEURT and ROUGE-L scores, and judgments from GPT-4 few-shot ICL and human evaluations. These examples are generated by GPT-4, as the underlying language model.
Metric/Method Reference Instance and Its Replica by Guided Instruction
@@
Reference Instance: Review:  Bromwell High is a cartoon comedy. It ran at the same time as some other programs about schoollife, such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’ssatire is much closer to reality than is ”Teachers”. The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ……… at ………. High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!Label: 1 (positive)
@@
BLEURT Score: 0.99 ROUGE-L Score: 1.00 GPT-4 ICL: Exact Match Human Eval.: Exact Match Instruction: You are provided with the first piece of an instance from the train split of the IMDB dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: 1 (positive) First Piece: Bromwell High is a cartoon comedy. It ran at the same time as some other programs about school life, such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’s satire is much closer to reality than is ”Teachers”. Second Piece: The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ......... at .......... High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!
@@
BLEURT Score: 0.87 ROUGE-L Score: 0.82 GPT-4 ICL: Near-exact Match Human Eval.: Near-exact Match Instruction: You are provided with Sentence 1 from the train split of the RTE dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: Although they were born on different planets, Oscar-winning actor Nicolas Cage’s new son and Superman have something in common - both were named Kal-el. Label: 1 (entailment) Sentence 2: Nicolas Cage’s new son is named Kal-el.
Reference Instance: Summary:  Kim is about to tell mom that Harry bought a new sofa, and he needs grey pillows.
BLEURT Score: 0.48 ROUGE-L Score: 0.12 GPT-4 ICL: Inexact Match Human Eval.: Inexact Match Instruction: You are provided with the first piece of a summary from the test split of the SAMSum dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: Kim is about to tell mom that Harry bought Second Piece: a new car but is worried mom will be upset. Kim is advised to tell mom in a positive way, focusing on Harry’s happiness.
@@
Metric Reference Instance and Its Replica by General Instruction
Reference Instance: Review:  Bromwell High is a cartoon comedy. It ran at the same time as some other programs about schoollife, such as ”Teachers”. My 35 years in the teaching profession lead me to believe that Bromwell High’ssatire is much closer to reality than is ”Teachers”. The scramble to survive financially, the insightful students who can see right through their pathetic teachers’ pomp, the pettiness of the whole situation, all remind me of the schools I knew and their students. When I saw the episode in which a student repeatedly tried to burn down the school, I immediately recalled ……… at ………. High. A classic line: INSPECTOR: I’m here to sack one of your teachers. STUDENT: Welcome to Bromwell High. I expect that many adults of my age think that Bromwell High is far fetched. What a pity that it isn’t!Label: 1 (positive)
@@
Reference Instance: Summary:  Kim is about to tell mom that Harry bought a new sofa, and he needs grey pillows.
@@
Appendix D Statistical Analysis: Bootstrap Resampling
We examine the statistical significance of results stemming from guided versus general instructions. Bootstrap resampling technique, involving 10,000 samples in the resampling process, is employed for this investigation (Efron 1979; Efron & Tibshirani 1993; Efron 2003). We concentrate on the alternative hypothesis that posits guided instructions produce outcomes closer to reference instances than those generated from general instructions, as evaluated by fluency, quality, and similarity. The performance metrics utilized here are BLEURT and ROUGE-L scores. We regard the ROUGE-L and BLEURT scores as statistically significant if the $p$-values $\leq 0.05$. We list all the computed $p$-values in Table 8 and highlight the statistically significant results by marking them with an asterisk in Table 4.
Table 8: $p$-values for differences between BLEURT and ROUGE-L scores of guided and general instructions, computed using bootstrap resampling with 10,000 resampling samples. $p$-values $\leq 0.05$ indicate statistically significant results.
Model Metric Split Instruction Datasets
IMDB AG News Yelp RTE WNLI SAMSum XSum
GPT-4 BLEURT Train Guided 0.319 0.005 0.981 0.041 0.000 0.478 0.115
Test/Valid Guided 1.000 0.000 1.000 0.075 0.035 0.283 0.170
ROUGE-L Train Guided 0.017 0.000 0.073 0.000 0.000 0.424 0.000
Test/Valid Guided 0.509 0.000 0.465 0.165 0.003 0.105 0.000
GPT-3.5 BLEURT Train Guided 1.000 0.006 1.000 0.465 0.008 0.746 0.093
Test/Valid Guided 0.992 0.134 0.932 0.030 0.020 0.293 0.321
ROUGE-L Train Guided 0.374 0.000 0.000 0.968 0.000 0.312 0.068
Test/Valid Guided 0.190 0.042 0.000 0.051 0.044 0.147 0.152
Appendix E Continued Training of GPT-3.5 Base Model for Intentional Contamination
For our validation study for contamination using the GPT-3.5 base model, we employ the previously referenced snapshot, gpt-3.5-turbo-0613. To conduct continued training on GPT-3.5, we submit a fine-tuning job via the OpenAI API. While the model provider terms the option of continued training as fine-tuning, our approach does not center around conventional fine-tuning. Our objective is to reproduce what the LLM—in our case, GPT-3.5—potentially observed during its pre-training phase when exposed to web data. To achieve this, we format the data in a way that encompasses the dataset title and its associated division, coupled with the entire details of the instance. We embed this information since it represents the minimal metadata an instance might possess when extracted from web data.
All data formats we used to introduce data contamination are listed in Table 9. Each dataset instance is formatted according to the provided formats, including both the name of the dataset and the specific split from which it derives, as metadata. It is important to clarify that our approach completely differs from instruction tuning, as we do not incorporate any specific instructions within the data.
Due to our project’s budget limitations and our emphasis on a manageable number of training samples, we opt to work with one dataset for each task in our validation study. In particular, we take 100 random samples, ensuring they were evenly distributed based on the label, from the training splits of the AG News, RTE, and XSum datasets to expose the GPT-3.5 base model. For training, all default hyperparameters set by OpenAI are maintained during our continued training phase. Upon training completion, we utilize particular checkpoints provided by OpenAI. For every experiment, the base model of GPT-3.5 is separately contaminated using each dataset split, resulting in three separate checkpoints, each associated with one of the aforementioned dataset splits.
Table 9: A complete list of all data formats used to contaminate the GPT-3.5 base model by further training. Each of these data formats is separately used to format every single instance with respect to the dataset task. Placeholders are as follows: {split_name} indicates the split name; {dataset_name} refers to the dataset name; {instance} represents a full instance in classification datasets; {sentence1} and {sentence2} stand for premise and hypothesis in NLI-based datasets; {document} and {summary} correspond to entire document and its summary for a single instance in the summarization datasets; and {label} is replaced with the input instance’s label where applicable.
Task Data Format
Classification This is an instance from the {split_name} split of the {dataset_name} dataset. Instance: {instance} Label: {label}
NLI This is an instance from the {split_name} split of the {dataset_name} dataset. Sentence 1: {sentence1} Sentence 2: {sentence2} Label: {label}
Summarization This is an instance from the {split_name} split of the {dataset_name} dataset. Document: {document} Summary: {summary}
Appendix F Examples of Replicas Generated Pre and Post Contamination of GPT-3.5
In Table 10, we showcase two examples of exact replicas derived from our controlled contamination study with GPT-3.5. These replicas are generated from the contaminated checkpoints obtained through additional training of the GPT-3.5 base model on the subset of the training partitions of the AG News and XSum datasets. Additionally, we highlight a near-exact match achieved from an instance in the training set of the GSM8k dataset when using GPT-4 as the underlying LLM. All these replicas are produced via the guided instructions.
Table 10: Examples of exact matches from GPT-3.5 after contamination with a subset of train splits of AG News and XSum datasets, alongside a near-exact instance reproduction from the GSM8k training set using GPT-4. Match types (exact/near-exact matches) are provided only for the generated completions from contaminated GPT-3.5 checkpoints, when contrasted with the completions from the GPT-3.5 base model (all completions by the GPT-3.5 base model are inexact matches). All completions are generated by guided instruction.
Method Reference Instance and Its Replica by Guided Instruction
Reference instance from the train split of AG News dataset: Article:  Report: Pennington Set to Sign Extension with Jets (Reuters) Reuters - The New York Jets andquarterback Chad Pennington are close to a deal that would make the quarterback the highest- paid player in franchise history.Label: 1 (Sports)
GPT-4 ICL: Exact Match Human Eval.: Exact Match Instruction: You are provided with the first piece of an instance from the train split of the AG News dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: 1 (Sports) First Piece: Report: Pennington Set to Sign Extension with Jets (Reuters) Reuters - The New York Jets and quarterback Chad Pennington are close to a deal that would make the Second Piece: ————————————————————————————————————————- GPT-3.5: 28-year-old signal caller the highest-paid player in franchise history, according to sources familiar with the negotiations. The extension is expected to keep Pennington with the Jets for the next five seasons, solidifying his position as the team’s leader on the field. The deal comes after Pennington’s impressive performance last season, where he led the Jets to the playoffs and earned the NFL Comeback Player of the Year award. With this extension, the Jets are showing their commitment to Pennington and their belief in his ability to lead the team to success. ————————————————————————————————————————- Contaminated GPT-3.5: quarterback the highest- paid player in franchise history.
Reference instance from the train split of XSum dataset: Summary:  Astronomers have found evidence for a planet being devoured by its star, yielding insights into the fate that will befall Earth in billions of years.
GPT-4 ICL: Exact Match Human Eval.: Exact Match Instruction: You are provided with the first piece of a one-sentence summary from the train split of the XSum dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: Astronomers have found evidence for a planet being devoured by its star, yielding insights Second Piece: ————————————————————————————————————————- GPT-3.5: into the process of planetary destruction and the fate of exoplanets. ————————————————————————————————————————- Contaminated GPT-3.5: into the fate that will befall Earth in billions of years.
Reference instance from the train split of GSM8k dataset: Question:  John writes 20 pages a day. How long will it take him to write 3 books that are 400 pages each?
GPT-4 ICL: Near-exact Match Human Eval.: Near-exact Match Instruction: You are provided with the first piece of a question from the train split of the GSM8k dataset. Finish the second piece of the question as exactly appeared in the dataset. Only rely on the original form of the question in the dataset to finish the second piece. First Piece: John writes 20 pages a day. How long will it take him to write 3 Second Piece: ————————————————————————————————————————- GPT-4: books of 400 pages each?
Appendix G Detailed Description of Datasets
@@
Recognizing Textual Entailment (RTE) Dataset. The Recognizing Textual Entailment (RTE) dataset originates from a succession of annual textual entailment challenges. These datasets were combined by the authors of the benchmark using data from four different editions: RTE1 (Dagan et al. 2005), RTE2 (Haim et al. 2006), RTE3 (Giampiccolo et al. 2007), and RTE5 (Bentivogli et al. 2009). The examples within these datasets were primarily formulated using text from news and Wikipedia sources. To maintain consistency, all these datasets were adapted into a two-class split. For those datasets that initially consisted of three classes, the categories of ”neutral” and ”contradiction” were combined to form a single class termed ”not entailment”. The RTE dataset combined has 2,490 examples for training, 277 examples for validation, and 3,000 examples for testing.
Winograd Natural Language Inference (WNLI) Dataset. The WNLI (Winograd Natural Language Inference) dataset is a benchmark for natural language understanding tasks, particularly for evaluating coreference resolution and pronoun disambiguation in context. The dataset is derived from the original Winograd Schema Challenge (Levesque et al. 2012) and contains sentence pairs where a pronoun needs to be resolved by determining whether it refers to the same entity as the previous sentence. While the dataset has a balanced training set between two classes, the test set is imbalanced, with 635 training examples, 146 testing examples, and 71 validation examples.
@@
Grade School Math 8k (GSM8k) Dataset. The GSM8k dataset is a curated dataset consisting of 8,500 linguistically diverse grade school math word problems, crafted meticulously by human authors. This collection is divided into 7,500 training examples and 1,000 designated for testing. The complexity of these problems varies, requiring between 2 to 8 sequential steps for resolution. Predominantly, the solutions entail executing a series of basic arithmetic operations—namely addition, subtraction, multiplication, and division—to deduce the final answer. This dataset is ideal for tasks involving multi-step mathematical reasoning.